# thanks: Corresponding author

### v2

@@ II.3.1 Example: $d=3$ and $\mu=1/2$
For $\mu=1/2$, and by choosing the nonincreasing order for probabilities, i.e., $q_{0}=q_{+}^{\mu}$, $q_{1}=q_{0}^{\mu}$, and $q_{2}=q_{-}^{\mu}$, we find from Eqs. (29) and (30)
@@
Using these in Eqs (19) and (24) to (26), we get $\eta_{0}=0$, $\eta_{1}=\xi_{0}=\xi_{1}=\pi$. Equation (28) gives $\theta^{\mu}=\pi$, which gives for the case (a)
@@
Using these in Eq. (18), one can find $\mathcal{U}$, from which $\mathcal{V}$ and $\mathcal{W}$ can also be obtained as
@@
Obviously, all three unitaries generate the same unistochastic matrix $\mathcal{Q}$. They are also mutually equi-biased as
@@
where generate three equivalent unistochastic matrices $\mathcal{Q}$, $\mathcal{Q}^{(123)}$, and $\mathcal{Q}^{(321)}$, respectively.
@@ III Entropic uncertainty relation for MEBs
For a two-qudit composite system $\mathcal{H}\otimes\mathcal{H}$, consider the following set of vectors
@@
where $\omega={\mathrm{e}}^{2\pi i/d}$, $k=1,\ldots,d-1$, and $\alpha=1,\ldots,L$. Here, $|e_{j}^{(\alpha)}\rangle^{\ast}$ denotes the complex conjugate of $|e_{j}^{(\alpha)}\rangle$ with respect to a definite basis, say the standard basis $\{|e_{j}^{(1)}\rangle\}_{j=0}^{d-1}$. Then
@@
for $k,k^{\prime}=1,\ldots,d-1$. Clearly, $N$ is anti-diagonal and its nonzero entries can be simplified as Eq. (4). In Appendix A, we derive the eigenvalues of $G$; in particular, its smallest and largest eigenvalues are $\lambda_{\min}(G)=1-(L-1)\lambda_{\max}(N)$ and $\lambda_{\max}(G)=1+(L-1)\lambda_{\max}(N)$, respectively. It follows therefore that the set of vectors (48) is linearly independent iff the Gram matrix is positive definite, i.e., $\lambda_{\min}(G)>0$, thereby completing the proof of part (a).
@@
By multiplying $\langle\phi_{l}^{(\beta)}|$ from the left on this equation, we get
@@
This implies that $CG^{\textrm{T}}={\mathbb{I}}_{L(d-1)}$, or equivalently, $C^{\textrm{T}}=G^{-1}$, where T stands for the transposition. Accordingly, up to a transposition, the matrix of the coefficients is obtained by the inverse of the Gram matrix $G$. This completes the proof of the lemma. ∎
@@ III.1 Inequality for the probabilities of projective measurements in MEBs of a qudit system
Building on the method presented in Ref. MolmerPRA2009 , we define a two-qudit pure state $\varrho\otimes{\mathbb{I}}]|\phi\rangle$, where $\varrho$ is an arbitrary quantum state on $\mathcal{H}$, and
@@
is a maximally entangled state of the composite system $\mathcal{H}\otimes\mathcal{H}$. One can easily show that $\langle\phi|\phi_{k}^{(\alpha)}\rangle=0$, for $k=1,\ldots,d-1$ and $\alpha=1,\ldots,L$. Invoking this and Lemma 4, vectors $\{|\phi\rangle,|\phi_{k}^{(\alpha)}\rangle\}$ span a $D$-dimensional subspace of $\mathcal{H}\otimes\mathcal{H}$ where D=L(d1)+1D=L(d-1)+1. Defining $\{|\chi_{l}\rangle\}_{l=1}^{d^{2}-D}$ as an orthonormal basis for the orthogonal complement of this subspace, and adding these vectors to the previously defined set of vectors we obtain, in general, a nonorthogonal basis for $\mathcal{H}\otimes\mathcal{H}$ as $\{|\phi\rangle,|\phi_{k}^{(\alpha)}\rangle,|\chi_{l}\rangle\}$ where $k=1,\ldots,d-1$ and $\alpha=1,\ldots,L$, and $l=1,\ldots,d^{2}-D$.
@@
where the first term follows from $\langle\phi|[\varrho\otimes{\mathbb{I}}]|\phi\rangle=1/d$, and the unnormalized vector $|\chi\rangle$ represents the component of $\varrho\otimes{\mathbb{I}}]|\phi\rangle$ that lies in the orthogonal subspace spanned by $\{|\chi_{l}\rangle\}$. For the second term, using Lemma 4, we have $\langle\widetilde{\phi}_{k}^{(\alpha)}|[\varrho\otimes{\mathbb{I}}]|\phi\rangle=[G^{-1}\Omega]_{k}^{(\alpha)}$, where $G^{-1}$ refers to the inverse of the Gram matrix $G$ defined by Eq. (III). The L(d1)L(d-1)-dimensional vector $\Omega$ is defined by its components $\Omega_{l}^{(\beta)}$, as
@@
where $\lambda_{\min}(G^{-1})$ is the smallest eigenvalue of $G^{-1}$, and the squared norm $\|\Omega\|^{2}=\Omega^{\dagger}\Omega$ can be expressed as
@@
where the last line follows from the fact that the Gram matrix of a linearly independent set of vectors is positive definite, as such $\lambda_{\max}(G)=1/\lambda_{\min}(G^{-1})$. To complete the proof, it remains only to determine the eigenvalues of the Gram matrix $G$. A complete derivation of the eigenvalues is provided in Appendix A—from which the maximum eigenvalue we need here is given by $\lambda_{\max}(G)=f_{L}(q)$, where $f_{L}(q)$ is defined by Eq. (56). This concludes the proof. ∎
In proving the theorem, we observe two different features of the d(L1)d(L-1)-dimensional vector $\Omega$, and the $(d-1)\times(d-1)$ matrix $N$; The vector $\Omega$ contains the probabilities of obtaining different outcomes when the system—initially prepared in state $\varrho$—is measured in MEBs, meanwhile, the Gram matrix $G$, so the submatrix $N$, provides all information on the MEBs. This last feature enables us to parameterize the probability distribution $q=(q_{0},\ldots,q_{d-1})$ associated with each MEBs, using d1d-1 independent parameters provided by the entries of the matrix $N$, as $N_{k,d-k}=\delta_{2k-1}+i\delta_{2k}$, where (d1)(d-1) real parameters $\delta_{1},\ldots,\delta_{d-1}$ are defined by Eqs. (5) and (6), respectively, (see Sec. II.1.2).
Quantum incompatibility imposes limitation on the information that can be extracted from measurements of different observables, and the uncertainty relation quantifies how the incompatibility of measurements restricts the precision with which we can simultaneously obtain information about incompatible observables. Entropic uncertainty relation provides a limit on how much information you can gain from measuring incompatible measurements. The inequality given by Theorem 5 for probabilities of projective measurement in MEBs can be used to construct an entropic uncertainty inequality.
@@ IV Entanglement detection
A powerful tool for detecting entanglement is the use of positive but not completely positive maps. A linear map $\Phi:\mathcal{B}(\mathcal{H}_{1})\rightarrow\mathcal{B}(\mathcal{H}_{2})$ is called positive if $\Phi X\geq 0\;\forall X\geq 0$, where $\mathcal{B}(\mathcal{H})$ denotes the space of bounded operators acting on the Hilbert space $\mathcal{H}$. The map $\Phi$ is called completely positive if and only if its trivial extension $\mathrm{id}_{n}\otimes\Phi$ also remains positive for all dimension $n$. For a given positive but not completely positive map $\Phi$, one can construct the corresponding entanglement witness $W=(d-1)(\textrm{id}_{n}\otimes\Phi)|\phi_{d}^{+}\rangle\langle\phi_{d}^{+}|$, where $|\phi_{d}^{+}\rangle=\frac{1}{\sqrt{d}}\sum_{i=0}^{d}|ii\rangle$. It turns out that $W$ has nonnegative expectation value on any separable state $\rho$, i.e. ${\textrm{Tr}}{\rho W}\geq 0$ for all separable states. Accordingly, a state for which ${\textrm{Tr}}{\rho W}<0$ is entangled. For a given set of $L$ MUBs, the authors of ChruscinskiPRA2018 have recently introduced the following positive and trace preserving map
@@
is positive and trace preserving. Here, $\gamma$ is given by
@@
where $q_{\max}=\max_{k}\{q_{0},\ldots,q_{d-1}\}$ and $f_{L}(q)$ is defined by Eq. (56). Also, $\mathcal{O}^{(b)}$ is a set of orthogonal rotation in $\mathbb{R}^{d}$ around the axis $\boldsymbol{n}_{\ast}=(1,1,\ldots,1)/\sqrt{d}$, so that $\mathcal{O}^{(b)}\boldsymbol{n}_{\ast}=\boldsymbol{n}_{\ast}$.
@@
First note that the map $\Phi$, defined by Eq. (7), is trace preserving, i.e., ${\textrm{Tr}}[\Phi X]={\textrm{Tr}}[X]$. To show that it is also positive, we require that $\Phi X\geq 0$ for all $X\geq 0$. According to the Mehta’s lemma, a Hermitian matrix $A$ is positive if ${\textrm{Tr}}[A^{2}]\leq{\textrm{Tr}}[A]/(d-1)$. In view of this, the map $\Phi$ is positive if and only if ${\textrm{Tr}}[\Phi\varrho]^{2}\leq{\textrm{Tr}}[\varrho]/(d-1)$ for all rank-one projection operators $P=|\psi\rangle\langle\psi|$. After some tedious but straightforward calculations one finds that
${\textrm{Tr}}[\Phi P]^{2}=\left[\frac{(d-1)^{2}-\gamma^{2}L^{2}}{d(d-1)^{2}}\right]+\frac{\gamma^{2}}{(d-1)^{2}}\textrm{T},$ (69)
where the expression T is defined by
@@
Remarkably, for a set of $L$ MUBs, the second term of the above expression reduces to L(L1)/dL(L-1)/d and the first term is bounded from above by 1+(L1)/d1+(L-1)/d, so that the condition ${\textrm{Tr}}[\Phi P]^{2}\leq 1/(d-1)$ is satisfied for $\gamma=1$. For MEBs, on the other hand, the second term is bounded from above by $L(L-1)q_{\max}$ where $q_{\max}=\max_{k}\{q_{k}\}$. For the fist term, however, the upper bound is given by
@@
Corresponding to the map $\Phi$, defined by Eq. (7), one can define the entanglement witness as follows
@@
where $0\leq\omega\leq 1$ and $|\phi_{d}^{+}\rangle=\frac{1}{\sqrt{d}}\sum_{i=0}^{d}|ii\rangle$. It is shown that isotropic states are entangled if and only if $\omega>1/(d+1)$. Straitforward calculations show that
@@
where in the the second equality we have used Eq. (2), and the third equality follows from the fact that the optimum detection happens when ${\textrm{Tr}}[\mathcal{O}^{(\alpha)}]$ takes its maximum value $d$, i.e., when $\mathcal{O}^{(\alpha)}={\mathbb{I}}_{d}$ for all $\alpha$. It turns out that $W$ can detect the entanglement of isotropic states when $\omega>\frac{1}{L\gamma}$. Accordingly, entanglement detection depends on the product of two parameters $L$ and $\gamma$: how many pairs of MEBs are used to detect entanglement, and how effective the measurements are. This provides a tradeoff relation between the number of measurements and their effectiveness in the sense that the more incompatible the measurements are, the fewer measurements are needed for detection. Clearly, complete entanglement detection is possible when a full set of $d+1$ measurements on MUBs are performed.
@@
Figure 5 illustrates this quantity-quality tradeoff in terms of the index $\mu$ for $d=3$ and different values of $L=2,3,4$. In this case, $\lambda_{\max}(N)=\sqrt{(3\mu-1)/2}$, as such, the constraint (49) leads to $\mu<(2/(L-1)^{2}+1)/3$. This yields the bounds $\mu<1$, $\mu<1/2$, and $\mu<11/27$, for $L=2$, $L=3$, and $L=4$, respectively. In all cases, the solid curves represent $(L\gamma)^{-1}$ when MEBs are used to construct witnesses. The dashed lines, on the other hand, capture the bounds when a set of $L$ MUBs are used to construct witnesses, so that $\gamma=1$. Not surprisingly, for a fixed $L$ the MUBs-based witness is finer than the MEBs-based witness, however, if a MEBs-based witness employs a larger set of measurements, it can surpass the MUBs-based witness in detecting entanglement, particularly within the specific index $\mu$ of the underlying distribution $q$ of the MEBs.
@@
Figure 5: Plots of $(L\gamma)^{-1}$ in terms of index $\mu$ for $d=3$ and different values of $L=2,3,4$. For equal $L$s, the MUBs-based witness (dashed lines) always captures a wider range of states relative to the MEBs-based witness (solid curves). However, if a MEBs-based witness employs a larger set of measurements, it can surpass the MUB-based witness in detecting entanglement, particularly within low values of the index $\mu$. As discussed in the main text, for $L=4$, the plot is restricted to $\mu\in[1/3,11/27)$, due to the imposed limit on $\mu$.
@@
The inequality $\lambda_{\max}(N)<1/(L-1)$ imposes a tradeoff relation on the MEBs, in the sense that for a given set of MEBs associated with a probability distribution $q=(q_{0},\ldots,q_{d-1})$, only $L<1/\lambda_{\max}(N)+1$ of them are relevant for entanglement detection, even if more than $L$ could be constructed. The constraint does not make sense for MUBs, as $N_{k,d-k}=0$ for the uniform distribution $q=(1/d,\ldots,1/d)$, so they are already optimal. For MEBs, however, it arises from the possibility of linear dependence among the maximally entangled states in Eq. (57), which renders certain bases redundant for the entanglement detection. In view of this, the significance of MEBs is pronounced when their distribution is sufficiently close to the uniform distribution. Despite the constraint imposed on the MEBs, they may still be valuable in certain dimensions, such as $d=6$, where the maximum number of $3$ MUBs has been identified. This raises the question that, is it possible for a set of at least $4$ MEBs to exist in dimension $d=6$? and, if such a set is found, can it be sufficiently close to the uniform distribution such that the constraint does not exclude any of the bases? This possibility merits further investigation.

### v3

@@ II.3.1 Example: $d=3$ and $\mu=1/2$
For $\mu=1/2$ and by choosing the nonincreasing order for probabilities, i.e., $q_{0}=q_{+}^{\mu}$, $q_{1}=q_{0}^{\mu}$, and $q_{2}=q_{-}^{\mu}$, we find from Eqs. (29) and (30)
@@
Using these in Eqs (19) and (24)-(26), we get $\eta_{0}=0$, $\eta_{1}=\xi_{0}=\xi_{1}=\pi$. Equation (28) gives $\theta^{\mu}=\pi$, which gives for case (a)
@@
Using these in Eq. (18), we can find $\mathcal{U}$, from which $\mathcal{V}$ and $\mathcal{W}$ can also be obtained as
@@
Obviously, all three unitaries generate the same unistochastic matrix $\mathcal{Q}$. They are also mutually equibiased as
@@
which generate three equivalent unistochastic matrices $\mathcal{Q}$, $\mathcal{Q}^{(123)}$, and $\mathcal{Q}^{(321)}$, respectively.
@@ III Entropic uncertainty relation for MEBs
For a two-qudit composite system $\mathcal{H}\otimes\mathcal{H}$, consider the set of vectors
@@
where $\omega={\mathrm{e}}^{2\pi i/d}$, $k=1,\ldots,d-1$, and $\alpha=1,\ldots,L$. Here $|e_{j}^{(\alpha)}\rangle^{\ast}$ denotes the complex conjugate of $|e_{j}^{(\alpha)}\rangle$ with respect to a definite basis, say the standard basis $\{|e_{j}^{(1)}\rangle\}_{j=0}^{d-1}$. Then
@@
for $k,k^{\prime}=1,\ldots,d-1$. Clearly, $N$ is anti-diagonal and its nonzero entries can be simplified as Eq. (4). In Appendix B we derive the eigenvalues of $G$; in particular, its smallest and largest eigenvalues are $\lambda_{\min}(G)=1-(L-1)\lambda_{\max}(N)$ and $\lambda_{\max}(G)=1+(L-1)\lambda_{\max}(N)$, respectively. It follows therefore that the set of vectors (40) is linearly independent if and only if the Gram matrix is positive definite, i.e., $\lambda_{\min}(G)>0$, thereby completing the proof of part (a).
@@
By multiplying $\langle\phi_{l}^{(\beta)}|$ from the left in this equation, we get
@@
This implies that $CG^{\textrm{T}}={\mathbb{I}}_{L(d-1)}$, or equivalently $C^{\textrm{T}}=G^{-1}$, where T stands for the transposition. Accordingly, up to a transposition, the matrix of the coefficients is obtained by the inverse of the Gram matrix $G$. This completes the proof of the lemma. ∎
@@ III.1 Inequality for the probabilities of projective measurements in MEBs of a qudit system
Building on the method presented in Ref. MolmerPRA2009 , we define a two-qudit pure state $\varrho\otimes{\mathbb{I}}]|\phi\rangle$, where $\varrho$ is an arbitrary quantum state on $\mathcal{H}$ and
@@
is a maximally entangled state of the composite system $\mathcal{H}\otimes\mathcal{H}$. One can easily show that $\langle\phi|\phi_{k}^{(\alpha)}\rangle=0$ for $k=1,\ldots,d-1$ and $\alpha=1,\ldots,L$. Invoking this and Lemma 3, vectors $\{|\phi\rangle,|\phi_{k}^{(\alpha)}\rangle\}$ span a $D$-dimensional subspace of $\mathcal{H}\otimes\mathcal{H}$ where D=L(d1)+1D=L(d-1)+1. Defining $\{|\chi_{l}\rangle\}_{l=1}^{d^{2}-D}$ as an orthonormal basis for the orthogonal complement of this subspace, and adding these vectors to the previously defined set of vectors we obtain, in general, a nonorthogonal basis for $\mathcal{H}\otimes\mathcal{H}$ as $\{|\phi\rangle,|\phi_{k}^{(\alpha)}\rangle,|\chi_{l}\rangle\}$ where $k=1,\ldots,d-1$; $\alpha=1,\ldots,L$; and $l=1,\ldots,d^{2}-D$.
@@
where the first term follows from $\langle\phi|[\varrho\otimes{\mathbb{I}}]|\phi\rangle=1/d$ and the unnormalized vector $|\chi\rangle$ represents the component of $\varrho\otimes{\mathbb{I}}]|\phi\rangle$ that lies in the orthogonal subspace spanned by $\{|\chi_{l}\rangle\}$. For the second term, using Lemma 3, we have $\langle\widetilde{\phi}_{k}^{(\alpha)}|[\varrho\otimes{\mathbb{I}}]|\phi\rangle=[G^{-1}\Omega]_{k}^{(\alpha)}$, where $G^{-1}$ refers to the inverse of the Gram matrix $G$ defined by Eq. (III). The L(d1)L(d-1)-dimensional vector $\Omega$ is defined by its components $\Omega_{l}^{(\beta)}$ as
@@
where $\lambda_{\min}(G^{-1})$ is the smallest eigenvalue of $G^{-1}$ and the squared norm $\|\Omega\|^{2}=\Omega^{\dagger}\Omega$ can be expressed as
@@
where the last line follows from the fact that the Gram matrix of a linearly independent set of vectors is positive definite, as such $\lambda_{\max}(G)=1/\lambda_{\min}(G^{-1})$. To complete the proof, it remains only to determine the eigenvalues of the Gram matrix $G$. A complete derivation of the eigenvalues is provided in Appendix B, from which the maximum eigenvalue we need here is given by $\lambda_{\max}(G)=f_{L}(q)$, where $f_{L}(q)$ is defined by Eq. (48). This concludes the proof. ∎
In proving the theorem, we observe two different features of the d(L1)d(L-1)-dimensional vector $\Omega$ and the $(d-1)\times(d-1)$ matrix $N$: The vector $\Omega$ contains the probabilities of obtaining different outcomes when the system, initially prepared in state $\varrho$, is measured in MEBs, while the Gram matrix $G$, and thus the submatrix $N$, provides all information on the MEBs. This last feature enables us to parameterize the probability distribution $q=(q_{0},\ldots,q_{d-1})$ associated with each MEB, using d1d-1 independent parameters provided by the entries of the matrix $N$, as $N_{k,d-k}=\delta_{2k-1}+i\delta_{2k}$, where (d1)(d-1) real parameters $\delta_{1},\ldots,\delta_{d-1}$ are defined by Eqs. (5) and (6), respectively (see Sec. II.1.2).
Quantum incompatibility imposes limitation on the information that can be extracted from measurements of different observables, and the uncertainty relation quantifies how the incompatibility of measurements restricts the precision with which we can simultaneously obtain information about incompatible observables. The entropic uncertainty relation provides a limit on how much information one can gain from measuring incompatible measurements. The inequality given by Theorem 4 for probabilities of projective measurement in MEBs can be used to construct an entropic uncertainty inequality.
@@ IV Entanglement detection
A powerful tool for detecting entanglement is the use of positive but not completely positive maps. A linear map $\Phi:\mathcal{B}(\mathcal{H}_{1})\rightarrow\mathcal{B}(\mathcal{H}_{2})$ is called positive if $\Phi X\geq 0\;\forall X\geq 0$, where $\mathcal{B}(\mathcal{H})$ denotes the space of bounded operators acting on the Hilbert space $\mathcal{H}$. The map $\Phi$ is called completely positive if and only if its trivial extension $\mathrm{id}_{n}\otimes\Phi$ also remains positive for all dimension $n$. For a given positive but not completely positive map $\Phi$, one can construct the corresponding entanglement witness $W=(d-1)(\textrm{id}_{n}\otimes\Phi)|\phi_{d}^{+}\rangle\langle\phi_{d}^{+}|$, where $|\phi_{d}^{+}\rangle=\frac{1}{\sqrt{d}}\sum_{i=0}^{d}|ii\rangle$. It turns out that $W$ has nonnegative expectation value on any separable state $\rho$, i.e. ${\textrm{Tr}}{\rho W}\geq 0$ for all separable states. Accordingly, a state for which ${\textrm{Tr}}{\rho W}<0$ is entangled. For a given set of $L$ MUBs, the authors of ChruscinskiPRA2018 introduced the positive and trace-preserving map
@@
is positive and trace preserving. Her $\gamma$ is given by
@@
where $q_{\max}=\max_{k}\{q_{0},\ldots,q_{d-1}\}$ and $f_{L}(q)$ is defined by Eq. (48). Also, $\mathcal{O}^{(b)}$ is a set of orthogonal rotation in $\mathbb{R}^{d}$ around the axis $\boldsymbol{n}_{\ast}=(1,1,\ldots,1)/\sqrt{d}$ so that $\mathcal{O}^{(b)}\boldsymbol{n}_{\ast}=\boldsymbol{n}_{\ast}$.
@@
First note that the map $\Phi$, defined by Eq. (6), is trace preserving, i.e., ${\textrm{Tr}}[\Phi X]={\textrm{Tr}}[X]$. To show that it is also positive, we require that $\Phi X\geq 0$ for all $X\geq 0$. According to Mehta’s lemma, a Hermitian matrix $A$ is positive if ${\textrm{Tr}}[A^{2}]\leq{\textrm{Tr}}[A]/(d-1)$. In view of this, the map $\Phi$ is positive if and only if ${\textrm{Tr}}[\Phi\varrho]^{2}\leq{\textrm{Tr}}[\varrho]/(d-1)$ for all rank-1 projection operators $P=|\psi\rangle\langle\psi|$. After some tedious but straightforward calculations one finds that
${\textrm{Tr}}[\Phi P]^{2}=\left[\frac{(d-1)^{2}-\gamma^{2}L^{2}}{d(d-1)^{2}}\right]+\frac{\gamma^{2}}{(d-1)^{2}}T,$ (61)
where the expression $T$ is defined as
@@
Remarkably, for a set of $L$ MUBs, the second term of the above expression reduces to L(L1)/dL(L-1)/d and the first term is bounded from above by 1+(L1)/d1+(L-1)/d so that the condition ${\textrm{Tr}}[\Phi P]^{2}\leq 1/(d-1)$ is satisfied for $\gamma=1$. For MEBs, on the other hand, the second term is bounded from above by $L(L-1)q_{\max}$ where $q_{\max}=\max_{k}\{q_{k}\}$. For the fist term, however, the upper bound is given by
@@
Corresponding to the map $\Phi$, defined by Eq. (6), one can define the entanglement witness as
@@
where $0\leq\omega\leq 1$ and $|\phi_{d}^{+}\rangle=\frac{1}{\sqrt{d}}\sum_{i=0}^{d}|ii\rangle$. It is shown that isotropic states are entangled if and only if $\omega>1/(d+1)$. Straightforward calculations show that
@@
where in the the second equality we have used Eq. (2), and the third equality follows from the fact that the optimum detection happens when ${\textrm{Tr}}[\mathcal{O}^{(\alpha)}]$ takes its maximum value $d$, i.e., when $\mathcal{O}^{(\alpha)}={\mathbb{I}}_{d}$ for all $\alpha$. It turns out that $W$ can detect the entanglement of isotropic states when $\omega>\frac{1}{L\gamma}$. Accordingly, entanglement detection depends on the product of two parameters $L$ and $\gamma$: how many pairs of MEBs are used to detect entanglement and how effective the measurements are. This provides a tradeoff relation between the number of measurements and their effectiveness in the sense that the more incompatible the measurements are, the fewer measurements are needed for detection. Clearly, complete entanglement detection is possible when a full set of $d+1$ measurements on MUBs is performed.
@@
Figure 3 illustrates this quantity-quality tradeoff in terms of the index $\mu$ for $d=3$ and different values of $L=2,3,4$. In this case, $\lambda_{\max}(N)=\sqrt{(3\mu-1)/2}$; as such, the constraint (41) leads to $\mu<(2/(L-1)^{2}+1)/3$. This yields the bounds $\mu<1$, $1/2$, and $11/27$, for $L=2$, 33, and $4$, respectively. In all cases, the solid curves represent $(L\gamma)^{-1}$ when MEBs are used to construct witnesses. The dashed lines, on the other hand, capture the bounds when a set of $L$ MUBs is used to construct witnesses so that $\gamma=1$. Not surprisingly, for a fixed $L$ the MUB-based witness is finer than the MEB-based witness; however, if a MEB-based witness employs a larger set of measurements, it can surpass the MUB-based witness in detecting entanglement, particularly within the specific index $\mu$ of the underlying distribution $q$ of the MEBs.
@@
Figure 3: Plot of $(L\gamma)^{-1}$ in terms of index $\mu$ for $d=3$ and different values of $L=2,3,4$. For equal $L$’s, the MUB-based witness (dashed lines) always captures a wider range of states relative to the MEB-based witness (solid curves). However, if a MEB-based witness employs a larger set of measurements, it can surpass the MUB-based witness in detecting entanglement, particularly within low values of the index $\mu$. As discussed in the text, for $L=4$, the plot is restricted to $\mu\in[1/3,11/27)$, due to the imposed limit on $\mu$.
@@
The inequality $\lambda_{\max}(N)<1/(L-1)$ imposes a tradeoff relation on the MEBs, in the sense that for a given set of MEBs associated with a probability distribution $q=(q_{0},\ldots,q_{d-1})$, only $L<1/\lambda_{\max}(N)+1$ of them are relevant for entanglement detection, even if more than $L$ could be constructed. The constraint does not make sense for MUBs, as $N_{k,d-k}=0$ for the uniform distribution $q=(1/d,\ldots,1/d)$, so they are already optimal. For MEBs, however, it arises from the possibility of linear dependence among the maximally entangled states in Eq. (49), which renders certain bases redundant for the entanglement detection. In view of this, the significance of MEBs is pronounced when their distribution is sufficiently close to the uniform distribution. Despite the constraint imposed on the MEBs, they may still be valuable in certain dimensions, such as $d=6$, where the maximum number of three MUBs has been identified. This raises the question of whether it is possible for a set of at least four MEBs to exist in dimension $d=6$ and, if such a set is found, can it be sufficiently close to the uniform distribution such that the constraint does not exclude any of the bases. This possibility merits further investigation.